OpenStudy (anonymous):

Identify which equation is a quadratic equation. Answer y = -5x - 3 y = -4x(x + 10) + 34 y + x^2 = (x - 4)(x + 3) y - 2 = (x - 2)(x^2 + 1)

OpenStudy (anonymous):

Quadratic equation rules: A quadratic equation is an equation of the form y = ax2 + bx + c, where “a”, “b” and “c” are real numbers, and “a” cannot be zero. y = ax2 + bx + c is called the Standard Form of a quadratic equation. A quadratic equation always has a degree of two. There must be an x2 term, and no other higher powers of x.

OpenStudy (anonymous):

So: A) Not in the form y = ax^2 + bx + c B) y = -4x(x + 10) + 34 is also y = -4x^2 -40x + 34 Do you have to go any further? B is the answer